OpenStudy (anonymous):

Prove each identity: a) cos (theta) + sin (theta) tan (theta) = 1/cos (theta) b) tan (theta) = 1-cos2(theta)/sin2(theta)

OpenStudy (solomonzelman):

cos (theta) + sin (theta) tan (theta) = 1/cos (theta) sin(theta) cos (theta) + sin (theta) --------- = 1/cos (theta) cos(theta) cos^2(theta) sin(theta) ---------- + sin (theta) --------- = 1/cos (theta) cos(theta) cos(theta) cos^2(theta) sin^2(theta) 1 ---------- + --------- = --------- cos(theta) cos(theta) cos(theta) cos^2(theta) +sin^2(theta) =1

OpenStudy (anonymous):

Thank you @SolomonZelman , can you help me with b as well?